### v1

# Statistics-encoded tensor network approach in disordered quantum many-body spin chains

Abstract
Simulating the dynamics of quantum many-body systems with disorder is a fundamental challenge. In this work, we propose a general approach—the statistics-encoded tensor network (SeTN)—to study such systems. By encoding disorder into an auxiliary layer and averaging separately, SeTN restores translational invariance, enabling a well-defined transfer matrix formulation. We derive a universal criterion, $n\gg\alpha^{2}t^{2}$, linking discretization $n$, disorder strength $\alpha$, and evolution duration $t$. This sets the resolution required for faithful disorder averaging and shows that encoding is most efficient in the weak-disorder, typically chaotic regime. Applied to the disordered transverse-field Ising model, SeTN shows that the spectral form factor is governed by the leading transfer-matrix eigenvalue, in contrast to the kicked Ising model. SeTN thus provides a novel framework for probing the disorder-driven dynamical phenomena in many-body quantum systems.
Introduction.—Simulating quantum many-body systems with disorder is crucial for understanding a wide range of phenomena, including quantum chaos [gutzwillerChaosClassicalQuantum1990a, dalessioQuantumChaosEigenstate2016b, haakeQuantumSignaturesChaos2018], many-body localization (MBL) [nandkishoreManyBodyLocalizationThermalization2015c, abaninColloquiumManybodyLocalization2019b], quantum spin glass [brayReplicaTheoryQuantum1980, sachdevGaplessSpinfluidGround1993], and criticality [fisherCriticalBehaviorRandom1995, pichCriticalBehaviorGriffithsMcCoy1998]. Recently, the chaotic dynamics of disordered systems predicted by random matrix theory (RMT) has attracted growing interest, driven by advances in experimental techniques [liEmergentUniversalQuench2024, dongMeasuringSpectralForm2025].
Analytical approaches to the disorder-averaged dynamics typically map them onto classical statistical models [chanSolutionMinimalModel2018c, bertiniExactSpectralForm2018, vonkeyserlingkOperatorHydrodynamicsOTOCs2018a, zhouEmergentStatisticalMechanics2019]. This mapping relies on Haar averaging [vonkeyserlingkOperatorHydrodynamicsOTOCs2018a, zhouEmergentStatisticalMechanics2019] or self-duality [chanSolutionMinimalModel2018c, bertiniExactSpectralForm2018] techniques and is primarily applicable to random circuits and Floquet systems. However, it cannot be directly applied to generic time-independent disordered Hamiltonians that usually lack self-duality. Numerical methods for studying disordered dynamics mainly include exact diagonalization (ED) and tensor network (TN) techniques. ED is limited to modest system sizes, e.g., $L\lesssim 24$ for the disordered XXZ chain [pietracaprinaShiftinvertDiagonalizationLarge2018a, sierantPolynomiallyFilteredExact2020a]. While TN methods can handle larger systems, they are mostly limited to the MBL regime [doggenManybodyLocalizationDelocalization2018a, chandaTimeDynamicsMatrix2020], as extending them to the ergodic phase remains challenging due to volume-law entanglement growth.
In this letter, we propose the statistics-encoded tensor network (SeTN), a general approach for studying disordered many-body systems with time-independent Hamiltonians. In SeTN, we encode disorder into an auxiliary layer and perform averaging separately (see Fig. 1(b)). By analyzing the decay rate of the singular values, we derive a criterion for the temporal resolution required for efficient disorder encoding: $n\gg\alpha^{2}t^{2}$, where $n$ is the number of time steps, $\alpha$ the disorder strength, and $t$ the evolution duration.
Our approach offers two key advantages: encoding is more efficient for weak disorder, a regime challenging for conventional tensor networks, and it restores spatial translation invariance, enabling a well-defined transfer matrix that allows studying dynamics under the thermodynamic limit and naturally integrates with Krylov iteration[stewartKrylovSchurAlgorithmLarge2002] and density matrix renormalization group (DMRG) methods [whiteDensityMatrixFormulation1992a, whiteDensitymatrixAlgorithmsQuantum1993]. As a demonstration, we investigate the spectral form factor (SFF) of the disordered transverse-field Ising model in its chaotic regime, revealing a prolonged pre-RMT transient dominated by a single, non-degenerate leading eigenvalue of the transfer matrix. This pathway to RMT behavior contrasts with Floquet models such as the kicked Ising model, where chaotic spectral statistics and eigenvalue degeneracies appear from the outset [chanSolutionMinimalModel2018c].
Statistics-encoded tensor network approach.—We consider a quantum many-body system with local disorder, described by the Hamiltonian $\mathcal{H}[\boldsymbol{h}]=H+\sum_{i=1}^{L}h_{i}H_{i}$, where $H$ is the disorder-free part, $\{H_{i}\}$ are local operators acting on site $i$, $\boldsymbol{h}=\{h_{i}\}$ is a disorder realization, and $L$ is the system size. Throughout this work we set $\hbar=1$. To simulate the real-time dynamics, we apply a first-order Trotter decomposition with a small time step $\tau\ll 1$, which gives the evolution operator $U(\tau)=\mathrm{e}^{-\mathrm{i}\tau H}\mathrm{e}^{-\mathrm{i}\tau\sum_{i}h_{i}H_{i}}+O(\tau^{2})$. This decomposition enables a TN representation of the full evolution operator (see Fig. 1(a)). For simplicity, we focus on the widely studied case where $H=\sum_{i}H_{i,i+1}$ includes only nearest-neighbor interactions and $H_{i}=\sigma^{z}_{i}$ is the Pauli-$z$ operator on site $i$. Generalizations to longer-range interactions or different disorder couplings are straightforward [supplementaryMaterial].
To proceed, as shown in Fig. 1(a), we first decompose the two-body unitary gates into two local tensors,
$\langle j_{n}^{[i]}j_{n+1}^{[i]}|\mathrm{e}^{-\mathrm{i}\tau H_{i,i+1}}|$ $j_{n}^{[i+1]}j_{n+1}^{[i+1]}\rangle=V^{(L)}_{j_{n}^{[i]}j_{n+1}^{[i]},k_{n}^{[i]}}V^{(R)}_{k_{n}^{[i]},j_{n}^{[i+1]}j_{n+1}^{[i+1]}},$ (1)
where $j_{n}^{[i]}$ refers to site $i$ and time step $n$, with repeated indices summed over implicitly. This can be achieved via singular value decomposition (SVD) or spectral decomposition. At the same time, we factor out the diagonal disorder term using a delta tensor: $\langle j_{n}^{[i]}|\mathrm{e}^{-\mathrm{i}\tau\sigma^{z}_{i}}|j_{n+1}^{[i]}\rangle=v_{l_{n}^{[i]}}\delta_{l_{n}^{[i]}j_{n}^{[i]}j_{n+1}^{[i]}}$, where $\delta_{ijk}$ is the three-order Kronecker tensor (equal to $1$ when $i=j=k$ and 0 otherwise), and $\boldsymbol{v}=[\mathrm{e}^{-\mathrm{i}\tau h_{i}},\mathrm{e}^{\mathrm{i}\tau h_{i}}]$ encodes the effect of local disorder. Throughout this work, we use the indices $j,k$ and $l$ to denote the time, space, and layer directions, respectively, as indicated by the coordinate system in Fig. 1. By combining these decompositions, we define a local tensor $W$ as
$W_{j_{n}^{[i]}j_{n+1}^{[i]},k_{n}^{[i-1]}k_{n}^{[i]},l_{n}^{[i]}}=\delta_{l_{n}^{[i]}j_{n}^{[i]}p_{n}^{[i]}}V^{(R)}_{k_{n}^{[i-1]},p_{n}^{[i]}q_{n}^{[i]}}V^{(L)}_{q_{n}^{[i]}j_{n+1}^{[i]},k_{n}^{[i]}}$, (2)
where $p$ and $q$ are auxiliary indices used for internal contractions. This construction is illustrated in the third column of Fig. 1(a), where the spatial slice at site $i$ corresponds to $\prod_{n}v_{l_{n}^{[i]}}W_{j_{n}^{[i]}j_{n+1}^{[i]},k_{n}^{[i-1]}k_{n}^{[i]},l_{n}^{[i]}}$.
Figure 1: (Color online) (a) Tensor network (TN) representation of the Trotterized evolvation operator. The tensor encoding the disorder is decomposed into a vector (blue circle) and a Kronecker delta tensor (black point). (b) A general two-layer TN representation of the disorder-averaged transfer matrix $\mathrm{T}$. The orange-shaded region indicates the disorder average. (c) Diagrammatic illustration of the SeTN compression. $S^{(n)}$ is the singular values obtained via corresponding singular value decomposition.
We study disorder-averaged dynamics, where many observables such as expectation values, the Rényi entropy, and the SFF admit a two-layer tensor-network representation, obtained by contracting the forward and backward evolutions $U$ and $U^{\dagger}$ (dark/light blue in Fig. 1(b)). The effect of disorder is encoded in a statistics layer (orange region), which after averaging renders all spatial sites equivalent. This naturally motivates the definition of a transfer matrix $\mathrm{T}$ that captures the combined action of unitary dynamics and disorder averaging across a single spatial slice,
$\mathrm{T}_{k_{1}^{\prime[i-1]},...,k_{n}^{\prime[i-1]},k_{1}^{[i-1]},...,k_{n}^{[i-1]};k_{1}^{\prime[i]},...,k_{n}^{\prime[i]},k_{1}^{[i]},...,k_{n}^{[i]}}$ (3)
$=\Big{\langle}\prod_{p=1}^{n}W^{*}_{j_{p}^{\prime[i]}j_{p+1}^{\prime[i]},k_{p}^{\prime[i-1]}k_{p}^{\prime[i]},l_{p}^{\prime[i]}}v^{*}_{l_{p}^{\prime[i]}}v_{l_{p}^{[i]}}W_{j_{p}^{[i]}j_{p+1}^{[i]},k_{p}^{[i-1]}k_{p}^{[i]},l_{p}^{[i]}}\Big{\rangle}$
$=\mathrm{O}_{l^{\prime}_{1},...,l^{\prime}_{n};l_{1},...,l_{n}}\prod W^{*}W,$
where the indices on the conjugated tensors are distinguished by primes and $\langle{\cdot}\rangle$ denotes the disorder average.
In the last line, We separate the disorder-averaged tensor $\mathrm{O}$ from the unitary layers (omitting site superscripts and tensor subscripts for clarity), which can be obtained explicitly by integrating over the distribution of $h$:
$\mathrm{O}_{l^{\prime}_{1},...,l^{\prime}_{n};l_{1},...,l_{n}}\equiv\Big{\langle}{\prod_{p=1}^{n}v_{l^{\prime}_{p}}^{*}v_{l_{p}}}\Big{\rangle}=\int\mathrm{d}h\;\mathrm{e}^{\mathrm{i}h\tau\sum_{p}(l^{\prime}_{p}-l_{p})}P(h)$ (4)
[…]
where $P(h)$ denotes the probability density function of the disorder variable $h$. We give explicit formulas for both uniform $\mathcal{U}(-\alpha,\alpha)$ and Gaussian $\mathcal{N}(0,\sigma^{2})$ disorder. While a similar expression for the Gaussian case was derived for KIM [bertiniExactSpectralForm2018], our result provides a more general formulation that applies to arbitrary systems with local disorder. This transfer matrix formulation provides a systematic and efficient framework for evaluating disorder-averaged quantities, particularly in the thermodynamic limit, and plays a central role in SeTN approach.
While the transfer matrix is a powerful tool, its implementation is limited by the exponential growth of the matrix $\mathrm{O}$, whose dimension $2^{n}\times 2^{n}$. We overcome this by a two-step SeTN compression shown in Fig. 1(c). First, each Se-Layer realization is written as a matrix product operator (MPO) with bond dimension 1; summing over $M$ such realizations via MPO addition [crosswhiteFiniteAutomataCaching2008] yields an MPO of bond dimension $M$ encoding the disorder average. Second, we iteratively compress this MPO using SVD. At $n=1$ (pink box in Fig. 1(c)), with singular values $S^{(1)}$ below a fixed threshold discarded. The truncated tensors are contracted with $A^{(2)}$ and decomposed again (blue box in Fig. 1(c)), and the process continues site by site to the final layer. Notely, since each step involves only local operations and the tensors $A^{(n)}$ are identically constructed via MPO addition from individual realizations, the overall memory requirement reduces to $O(M)$ [supplementaryMaterial], compared to the naive $O(nM^{2})$ scaling. This significantly enhances computational efficiency and enables averaging over a large number of realizations ($M\gg 1$) in practice.
Figure 2: (Color online) (a) Blue dots: squared relative singular values from SeTN compression under various parameters (Num.). Dashed lines: analytic predictions from perturbation theory (Ana.), with coefficients fitted to numerical data. (b) Blue dots: first seven coefficients $c_{i}$ obtained by fitting the data in (a). These are further fitted to exponential (green dashed), Gaussian (purple dashed), and gamma-like (red dot-dashed) functions. (c) Spectral form factor of the disordered TFIM with system size $L=4$, computed using four methods: exact diagonalizing 1000 disorder realizations (avg #1000, green dotted), numerical integration (Exact, blue dashed), SeTN (red solid), and perturbative results (Pert., gray dot-dashed). The mean deviation between SeTN and the exact result is approximately 0.021. Parameters: $J=b=1$, Trotter step $\tau=0.005$, averaging over $M=10^{6}$ realizations; singular values below $10^{-10}$ are truncated.
Properties of Se-Layer and SeTN effectiveness.—The previous section outlined the implementation of SeTN. A natural question is whether disorder averaging can be faithfully encoded into the TN representation. This effectiveness crucially depends on the behavior of the singular value spectrum $S^{(n)}$, illustrated in Fig. 1(b): only when the singular values decay rapidly, ideally exponentially, can the SeTN accurately compress the averaged transfer matrix with a low bond dimension. To analyze this, we derive an analytic expression for the spectrum by solving the recursive structure in SeTN compression perturbatively [supplementaryMaterial]:
$\left(S_{i}(t)/S_{0}(t)\right)^{2}\simeq c_{i}\left(\alpha^{2}\tau t\right)^{i},$ (5)
where $S_{i}(t)$ denotes the $i$-th singular value at the $n$-th compression step (with $n=t/\tau$), ordered in descending magnitude, $S_{0}(t)$ is the largest and $\alpha$ is the disorder strength. The coefficients $c_{i}$ depend only on the disorder statistics; for a uniform distribution, the first two values can be computed exactly as $c_{1}=2/3$ and $c_{2}=8/45$ [supplementaryMaterial]. This result shows that the decay of the singular values is governed by the dimensionless parameter $\alpha^{2}\tau t$, which confirms the exponential suppression of subleading components in the singular value spectrum when $\alpha^{2}\tau t\ll 1$, or equivalently,
$n\gg\alpha^{2}t^{2}.$ (6)
This condition ensures that SeTN achieves accurate, low-bond-dimension compression of disorder-averaged transfer matrices.
To validate this analytic result, we numerically investigate the singular value spectra under various parameter settings [supplementaryMaterial] and compare them with the perturbative result. As shown in Fig. 2(a), the blue dots represent the normalized squared singular values $(S_{i}(t)/S_{0}(t))^{2}$ obtained from numerical simulations using a uniform disorder distribution, while the red dashed lines show the corresponding analytic estimates. The excellent agreement between the numerical data and theoretical predictions in the region $\alpha^{2}\tau t\ll 1$ confirms the exponential decay behavior discussed above.
In Fig. 2(b), we further analyze the coefficients $c_{i}$ appearing in Eq. 6. The plot shows the first seven coefficients extracted by fitting the numerical spectra in Fig. 2(a) to the analytic form. Notably, these coefficients decay faster than exponentially but slower than Gaussian, and are well approximated by a gamma-like function, which highlights the intrinsic efficiency of the SeTN method. Importantly, despite the rapid decay, the leading coefficients remain large. For a typical setting where $\alpha^{2}\tau t=10^{-2}$ (e.g., $\alpha=0.1$, $\tau=0.01$, $t=100$), the corresponding contribution to $(S_{i}/S_{0})^{2}$ at $i=7$ is on the order of $10^{-14}$, several orders of magnitude smaller than the coefficient $c_{7}=10^{-6}$. Thus, the overall compression efficiency of SeTN is still primarily governed by the dimensionless combination $\alpha^{2}\tau t$, as emphasized earlier.
For fixed disorder strength $\alpha$ and time $t$, we provide a physical interpretation of the above properties: while increasing the discretization number $n$ leads to exponential growth in the dimension of the transfer matrix, the total information introduced by disorder averaging, quantified by the dimensionless parameter $\alpha^{2}t^{2}$, remains unchanged. In the SeTN approach, this information is distributed across a tensor network of length $n$, such that the amount of information encoded per time step scales as $\alpha^{2}t^{2}/n=\alpha^{2}\tau t$. This local information density is directly related to the decay of the singular value spectrum, which allows efficient compression and accurate representation of disorder-averaged dynamics when $n\gg\alpha^{2}t^{2}$.
Scaling behavior of spectral form factor.—The preceding analysis shows that SeTN is more efficient for weak disorder, as singular values decay faster. In this regime, the system is typically driven into chaos, which is challenging for conventional tensor networks. To illustrate SeTN’s ability to handle chaotic dynamics, we investigate the spectral form factor (SFF) of the disordered transverse-field Ising model (disordered TFIM) in its chaotic regime.
The SFF, defined as the Fourier transform of the two-point correlation function of the spectral density [supplementaryMaterial],
$K(t)=\langle{|\operatorname{Tr}U(t)|^{2}}\rangle=\operatorname{Tr}\langle{U(t)\otimes U^{*}(t)}\rangle,$ (7)
where $U(t)=\mathrm{e}^{-\mathrm{i}Ht}$ is the evolution operator, serves as one of the simplest analytically tractable probes of quantum chaos. It captures both short- and long-range spectral correlations and has been extensively used in many-body chaos studies [chanSolutionMinimalModel2018c, kosManyBodyQuantumChaos2018b, chanSpectralStatisticsSpatially2018c, bertiniExactSpectralForm2018, dongMeasuringSpectralForm2025], where it typically exhibits an initial transient followed by the ramp–plateau structure predicted by RMT [haakeQuantumSignaturesChaos2018, liuSpectralFormFactors2018b].
Here, we study the SFF in the disordered TFIM whose Hamiltonian reads, $\mathcal{H}[\boldsymbol{h}]=J\sum_{i=1}^{L-1}\sigma^{z}_{i}\sigma^{z}_{i+1}+b\sum_{i}^{L}\sigma^{x}_{i}+\sum_{i}h_{i}\sigma^{z}_{i}$, where $J$ is the coupling strength, $b$ is the magnetic field, and $h_{i}$ are independent random variables drawn from a uniform distribution in the range $\alpha,\alpha]$. Setting $J=b=1$ and $\alpha=0.5$ where the system is chaotic [supplementaryMaterial], we compute the SFF via ED for $L=9$ to $16$ averaging $2440$ independent disorder realizations for each size. The results clearly display the onset of the linear ramp as shown in Fig. 3(a).
Since the SFF is non-self-averaging [prangeSpectralFormFactor1997b], reliable evaluation typically requires extensive disorder averaging. By expressing the SFF into a two-layer tensor-network form, however, SeTN can be directly employed as outlined in the previous section. Fig. 2(c) shows that SeTN (red line) reproduces converged results obtained from numerical integration (blue dashed line) even at long times. For comparison, we include disorder-averaged ED ($10^{3}$ samples, green dotted) and second-order perturbation results (SeTN with bond dimension three [supplementaryMaterial], gray dash-dotted). The slight ED deviation (see inset of Fig. 2(c)) highlights SeTN’s accuracy, whereas the breakdown of perturbation theory emphasizes the importance of small singular values for long-time dynamics.
A key advantage of SeTN is its ability to restore spatial translational invariance broken by disorder, thereby enabling a well-defined transfer matrix and access to the thermodynamic limit. In Fig. 3(b), we show the largest eigenvalues of the transfer matrix, computed by Krylov iteration (blue crosses) and non-Hermitian DMRG (red circles) [supplementaryMaterial]. For reference, we fit the finite-size ED data in Fig. 3(a) to the form $K(t)=\lambda(t)^{L}$, with the extracted $\lambda(t)$ plotted as black lines. The excellent agreement between the largest eigenvalues and $\lambda(t)$ demonstrates that, within the accessible time window, the SFF is indeed governed by the non-degenerate leading eigenvalues of the transfer matrix. This behavior contrasts with Floquet models such as the kicked Ising model [chanSolutionMinimalModel2018c], where the SFF is controlled by a highly degenerate set of leading eigenvalues from the outset.
Within the accessible time window the dynamics are governed by the non-degenerate leading eigenvalue, which is size-independent, cannot explain the size-dependent onset of the RMT ramp (the Thouless time, black dashed in Fig. 3(a)). This raises the question of how the system crosses over from the leading-eigenvalue regime to the ramp regime. Direct evaluation of the eigenvalues at the Thouless time is hindered by rapid bond growth in Krylov iteration and the lack of convergence in non-Hermitian DMRG. Inspired by the observed near degeneracy of the leading eigenvalues where the confidence interval (shaded region) expands, we conjecture that around the Thouless time, the leading eigenvalues become nearly degenerate, with the degree of near-degeneracy increasing in time. Such increasing near degeneracy implies that at finite sizes, multiple eigenvalues contribute comparably, producing the ramp, whereas in the thermodynamic limit the single largest eigenvalue eventually dominates. This scenario naturally explains both the emergence of the RMT ramp and the size-dependence of the Thouless time, as further illustrated by a toy-model analysis [supplementaryMaterial]. These insights highlight the potential of SeTN-defined transfer matrices in capturing the crossover to RMT spectral statistics in disordered time-independent quantum systems.
Figure 3: (Color online) (a) SFF of the disordered TFIM with sizes $L=9$ to $16$ (top to bottom), computed by ED using 2440 independent disorder realizations for each size. (b) Magnitudes of the leading and subleading eigenvalues of the SeTN-derived transfer matrix as functions of time, obtained using the Krylov method (blue crosses) and non-Hermitian DMRG (red circles). Black lines represent the fitted coefficient extracted by fitting the data in panel (a) to the form $\lambda(t)^{L}$, and the shaded region indicates the 80% confidence interval. Parameters: Trotter step $\tau=0.05$, $J=b=1,\alpha=0.5$.
Summary and outlook.—We propose SeTN, a tensor network approach for studying chaotic dynamics in general disordered quantum systems. By encoding disorder into an auxiliary layer and performing the disorder averaging separately, SeTN restores spatial translational invariance and enables accurate simulations directly in the thermodynamic limit. From the singular value spectrum of SeTN, we derive and numerically confirm a criterion for efficient disorder encoding, $n\gg\alpha^{2}t^{2}$, which indicates that encoding is more efficient under weak disorder, a regime that is typically chaotic and challenging for conventional TN methods. We compute the leading eigenvalue of the SeTN-derived transfer matrix for the SFF in the disordered TFIM, and show that it accurately captures the scaling behavior of the SFF within the accessible time range. Based on the behavior within this range, we conjecture how the single-eigenvalue dominance crosses over to random-matrix-theory behavior, a key aspect of many-body quantum chaos that deserves further investigation.
This SeTN approach can be readily extended to other disordered models and physical observables that can be represented by multilayer TN representations. Examples include Rényi entropies [chanSolutionMinimalModel2018c, zhouEmergentStatisticalMechanics2019], out-of-time-ordered correlators (OTOCs)[nahumOperatorSpreadingRandom2018c, vonkeyserlingkOperatorHydrodynamicsOTOCs2018a, chanSolutionMinimalModel2018c], and spin-glass order parameters[huseLocalizationprotectedQuantumOrder2013a, pekkerHilbertGlassTransitionNew2014a]. The simple and general structure of SeTN is reminiscent of an effective trajectory pairing mechanism in the Feynman-trajectory picture [vonkeyserlingkOperatorHydrodynamicsOTOCs2018a, zhouEmergentStatisticalMechanics2019, zhouEntanglementMembraneChaotic2020], where disorder averaging cancels most quantum trajectories through destructive interference. SeTN may also serve as a bridge between local averaging schemes and more advanced concepts, such as influence functionals [nguyenCorrelationFunctionsTensor2024, sonnerInfluenceFunctionalManybody2021a, yeConstructingTensorNetwork2021, leroseInfluenceMatrixApproach2021b] and entanglement barriers [rathEntanglementBarrierIts2023], and offers a powerful platform for investigating the impact of disorder on dynamical phenomena such as diffusion and hydrodynamics [khemaniOperatorSpreadingEmergence2018c, rakovszkyDiffusiveHydrodynamicsOutofTimeOrdered2018c]. Overall, SeTN provides a unifying framework for exploring the complex interplay between locality, unitarity, disorder, and chaos in time-independent many-body quantum systems.
Acknowledgment.— This work was supported in part by National Natural Science Foundation of China Grants No. 12474353 and Beijing Natural Science Foundation (Grant No. 1232025). SJR acknowledges supports from Peng Huanwu Visiting Professor Program, Chinese Academy of Sciences, and Academy for Multidisciplinary Studies, Capital Normal University. D.W. acknowledges support from the National Research Foundation, Singapore and the Agency for Science, Technology and Research (A*STAR) under the Quantum Engineering Programme (NRF2021-QEP2-02-P03) and from CQT Core Funding Grant CQT_SUTD_2025_01.
Supplemental Material:
Statistics-encoded tensor network approach in disordered quantum many-body spin systems
I Generalization of the SeTN Framework
In the main text, we focus on a simplified case where the clean Hamiltonian consists of nearest-neighbor interactions, i.e., $H=\sum_{i}H_{i,i+1}$ and $H_{i}=\sigma^{z}_{i}$, with $\sigma^{z}_{i}$ being the Pauli-$z$ operator at site $i$. Here, we outline how the statistics-encoded tensor network (SeTN) framework can be generalized to broader classes of Hamiltonians.
When $H_{\text{clean}}$ contains interactions beyond nearest neighbors, the corresponding step evolution operator can still be represented by matrix product operators (MPOs). This is enabled by automata-based constructions [crosswhiteFiniteAutomataCaching2008] or methods based on the time-dependent variational principle (TDVP) [haegemanUnifyingTimeEvolution2016a]. Specifically, we first express disorder-free term $H$ in MPO form. By acting the evolution operator on the identity matrix (or equivalently, the infinite-temperature state), i.e., $\mathrm{e}^{-\mathrm{i}\tau H}\cdot I$, we obtain an MPO representation of the evolution operator. This structure is illustrated in the first row of Fig. 4, where we denote the MPO tensor at site $n$ by $B^{(n)}$.
Figure 4: Tensor network representation of the generalized SeTN framework. (Left) Clean Hamiltonian evolution represented as an MPO with local tensors $B^{(n)}$. (Right) Representation of a general local disorder term $e^{-\mathrm{i}\tau h_{i}H_{i}}$ via eigen-decomposition. The combination of these elements modifies the local tensor $W$ in the SeTN construction while preserving its overall structure.
For general local operators $H_{i}$ that are not diagonal in the computational basis, we can still apply the SeTN framework by performing an eigen-decomposition:
$(\mathrm{e}^{-\mathrm{i}\tau h_{i}H_{i}})_{jp}=u_{jq}v_{q}\delta_{qr}u_{rp}^{*}=v_{l}\big{(}u_{jq}u_{rp}^{*}\delta_{lqr}\big{)},$
as shown on the right side of Fig. 4. Here, we omit the layer index $n$ and site label $(i)$ for clarity, and summation over repeated indices is implied.
With these preparations, the local tensor $W$ used in the main text is modified as follows:
$W_{j_{n}^{[i]}j_{n+1}^{[i]},k_{n}^{[i-1]}k_{n}^{[i]},l_{n}^{[i]}}=u_{j_{n}^{[i]}q_{n}^{[i]}}\delta_{l_{n}^{[i]}q_{n}^{[i]}r_{n}^{[i]}}u_{r_{n}^{[i]}p_{n}^{[i]}}^{*}B^{(i)}_{p_{n}^{[i]}j_{n+1}^{[i]},k_{n}^{[i-1]}k_{n}^{[i]}}$.
Despite the more general setting, the tensor product structure at site $i$ remains unchanged and takes the form:
$\prod_{n}v_{l_{n}^{[i]}}W_{j_{n}^{[i]}j_{n+1}^{[i]},k_{n}^{[i-1]}k_{n}^{[i]},l_{n}^{[i]}},$
as illustrated in the third column of Fig. 1(b) in the main text.
This extension allows the SeTN framework to accommodate arbitrary local disorder terms, including those with off-diagonal components, while retaining the key structure that facilitates efficient tensor network computations.
II Memory efficiency in SeTN truncation
In the SeTN truncation method described in the main text, one naively needs to store $n$ tensors $A^{(1)},\cdots,A^{(n)}$, each of dimension $D^{2}M^{2}$, where $D$ is the dimension of the internal (time) index $k$. This results in a total memory cost of $O(nD^{2}M^{2})$, which can be prohibitively large in practice.
This cost can be significantly reduced by exploiting the time-translation invariance of the system. After performing the MPO addition over $M$ realizations, all the tensors $A^{(n)}$ become identical for $2\leq n\leq N-1$, except for the two boundary layers. As a result, the required memory is reduced to $O(D^{2}M^{2})$.
Moreover, the structure of $A^{(n)}$ allows for further simplification. For a single realization ($M=1$), the tensor takes the form: $A^{(n)}_{l^{\prime}_{n}l_{n},11}=v_{l^{\prime}_{n}}^{*}v_{l_{n}}$ while for $M>1$, corresponding to $M$ independent realizations, we have:
$A^{(n)}_{l^{\prime}_{n}l_{n},j_{n-1}j_{n}}=v_{l^{\prime}_{n}}^{*(j_{n})}v_{l_{n}}^{(j_{n})}\delta_{j_{n-1},j_{n}},$
where $v_{l_{n}}^{(j_{n})}$ denotes the vector associated with the $j_{n}$-th realization.
Using this structure, the contraction in the iterative truncation procedure becomes:
$S^{(n-1)}_{p_{n-1}}V^{(n-1)}_{p_{n-1}j_{n-1}}A^{(n)}_{l^{\prime}_{n}l_{n},j_{n-1}j_{n}}=S^{(n-1)}_{p_{n-1}}V^{(n-1)}_{p_{n-1}j_{n-1}}v_{l^{\prime}_{n}}^{*(j_{n})}v_{l_{n}}^{(j_{n})}\delta_{j_{n-1},j_{n}}=\sum_{j_{n},p_{n-1}}S^{(n-1)}_{p_{n-1}}V^{(n-1)}_{p_{n-1}j_{n}}v_{l^{\prime}_{n}}^{*(j_{n})}v_{l_{n}}^{(j_{n})}.$
In the last equation, we have explicitly written out the summation indices for clarity.
In this expression, $S^{(n-1)}$ is a vector of size $\chi_{n-1}$, $V^{(n-1)}$ is a matrix of size $\chi_{n-1}\times M$, and both $v^{(j_{n})}$ vectors have dimension $D\times M$. The resulting tensor after summation has dimension $\chi_{n}\times M$, and the total memory cost of this contraction step is $O((\chi+D)M)$.
Therefore, the total memory complexity of the SeTN scheme is only $O(M)$, rather than $O(nD^{2}M^{2})$, which significantly improves computational efficiency.
III Analysis of the singular value spectrum in SeTN
IV.1 Perturbative analysis
This appendix presents a perturbative treatment of the MPO approximation depicted in Fig.  1 (d). The MPO is first reshaped into a matrix product state (MPS) with physical dimension $4$, upon which the approximation procedure can be formulated. The procedure is summarized diagrammatically as follows:
In this diagram, the circle represents a local vector $v_{k}(h)$, defined by
$v_{k}=\begin{cases}1,&k=1,4,\\ \mathrm{e}^{-2\mathrm{i}h\tau},&k=2,\\ \mathrm{e}^{2\mathrm{i}h\tau},&k=3,\end{cases}$
while the triangles (or rectangle) correspond to local matrices $A_{k}^{(n)}$, which is (not) isometric.
To be more precise, we aim as find a set of $\{A_{k}^{(n)}\}$ such that for any sequence $\{k_{1},k_{2},\dots,k_{N}\}$ with $k_{i}\in\{1,2,3,4\}$, we have,
$\int\mathrm{d}h\,P(h)v_{k_{1}}v_{k_{2}}\cdots v_{k_{N}}=\sum_{j_{1},j_{2},\cdots,j_{N+1}=1}^{m}\left(A_{k_{1}}^{(1)}\right)_{j_{1}j_{2}}\left(A_{k_{2}}^{(2)}\right)_{j_{2}j_{3}}\cdots\left(A_{k_{N}}^{(N)}\right)_{j_{N}j_{N+1}}.$
The random variable $h$ follows a probability distribution function $P(h)$, taken to be uniform on the interval $\alpha,\alpha]$, that is, $P(h)=\frac{1}{2\alpha}\left(\Theta(h+\alpha)-\Theta(h-\alpha)\right)$ where $\Theta(h)$ denotes the Heaviside step function.
To this end, we explicitly construct the local matrices $A_{k}^{(n)}$ and verify that they reproduce the integral structure up to leading perturbative corrections. The matrices $A_{k}^{(n)}$ are found to take the forms
$A_{1,4}^{(n)}$ $=\left[\begin{array}[]{ccc}\frac{1}{2}+\delta_{1}&0&\left(-\frac{1}{2}+\delta_{2}\right)\frac{1}{\sqrt{n(2n-1)}}\\ 0&\left(\frac{1}{2}+\frac{3}{10}\alpha^{2}\tau^{2}\right)\sqrt{\frac{n-1}{n}}&0\\ \frac{2\alpha^{4}\tau^{4}}{45}\sqrt{(n-1)(2n-3)}&0&\frac{1}{2}\sqrt{\frac{(n-1)(2n-3)}{n(2n-1)}}\end{array}\right],$ (8)
$A_{2,3}^{(n)}$ $=\left[\begin{array}[]{ccc}\frac{1}{2}-\delta_{1}&\pm\frac{1}{\sqrt{2n}}&\left(\frac{1}{2}+\delta_{2}\right)\frac{1}{\sqrt{n(2n-1)}}\\ \mp 4\delta_{1}\sqrt{\frac{n-1}{2}}&\left(\frac{1}{2}-\frac{3}{10}\alpha^{2}\tau^{2}\right)\sqrt{\frac{n-1}{n}}&\pm\sqrt{\frac{2(n-1)}{n(2n-1)}}\\ \frac{2\alpha^{4}\tau^{4}}{15}\sqrt{(n-1)(2n-3)}&\mp\frac{4\alpha^{2}\tau^{2}}{15}\sqrt{\frac{(n-1)(2n-3)}{2n}}&\frac{1}{2}\sqrt{\frac{(n-1)(2n-3)}{n(2n-1)}}\end{array}\right],$
where the perturbative corrections $\delta_{1}$ and $\delta_{2}$ are given by
$\delta_{1}$ $=\frac{\alpha^{2}\tau^{2}}{6}+\frac{(3-4n)\alpha^{2}\tau^{2}}{45},$
$\delta_{2}$ $=\frac{(4n-3)\alpha^{2}\tau^{2}}{6}-\frac{(24n^{2}-44n+21)\alpha^{4}\tau^{4}}{45}.$
The derivation proceeds in three main steps:
Step 1: Approximate $A^{(1)}$ by performing an eigenvalue decomposition of the averaged density matrix at the first ‘site’,
$|,\displaystyle A^{(1)}=\text{eig}\int\mathrm{d}h\,P(h)\ket{v}\hskip-1.99168pt\bra{v}$,
as illustrated in the following diagram:
Here the dashed line denotes a bond of dimension $1$, and the diamond shape represents the eigenvalues.
The integral is evaluated by expanding the vector $v$ in powers of $\tau$ and applying standard second-order perturbation theory. The resulting eigenvalues are,
$E^{(1)}=\left\{4-\frac{8\alpha^{2}\tau^{2}}{3}+\frac{16\alpha^{4}\tau^{4}}{9},\frac{8\alpha^{2}\tau^{2}}{3}-\frac{32\alpha^{4}\tau^{4}}{15},\frac{16\alpha^{4}\tau^{4}}{45},0\right\},$
and the corresponding eigenvectors are given by,
$A^{(1)}_{1,4}$= $\left[\begin{array}[]{ccc}\frac{1}{2}+\frac{\alpha^{2}\tau^{2}}{6}-\frac{\alpha^{4}\tau^{4}}{45}&0&-\frac{1}{2}+\frac{\alpha^{2}\tau^{2}}{6}-\frac{\alpha^{4}\tau^{4}}{45}\\ \end{array}\right],$
$A^{(1)}_{2,3}$= $\left[\begin{array}[]{ccc}\frac{1}{2}-\frac{\alpha^{2}\tau^{2}}{6}+\frac{\alpha^{4}\tau^{4}}{45}&\pm\frac{1}{\sqrt{2}}&\frac{1}{2}+\frac{\alpha^{2}\tau^{2}}{6}-\frac{\alpha^{4}\tau^{4}}{45}\end{array}\right].$
These expressions agree with the form shown in Eq. 8 at $n=1$, and serve as the initial condition for the recursion.
Step 2: Approximate $A^{(2)}$ based on the previously obtained $A^{(1)}$. To construct the density matrix required for eigenvalue decomposition, we introduce the renormalized tensor $w^{(2)}$, defined as,
$w^{(2)}_{i}$ $=\sum_{k}v_{k}\left(A^{(1)}_{k}\right)_{1,i}^{*}.$ (9)
The matrix $A^{(2)}$ is then obtained by performing an eigenvalue decomposition of the averaged outer product,
$A^{(2)}$ $=\text{eig}\int\mathrm{d}h\,P(h)|w^{(2)}\otimes v\rangle\hskip-1.42262pt\langle w^{(2)}\otimes v|,$
illustrated diagrammatically as,
As in Step 1, the integral is evaluated perturbatively to order $O(\tau^{4})$. The resulting eigenvalues are
$E^{(2)}=$ $\left\{16-\frac{64\alpha^{2}\tau^{2}}{3}+\frac{1616\alpha^{4}\tau^{4}}{45},\frac{64\alpha^{2}\tau^{2}}{3}-\frac{128\alpha^{4}\tau^{4}}{3},\frac{128\alpha^{4}\tau^{4}}{15},0\times 9\right\},$
where $0\times 9$ indicates ninefold degenerate zero eigenvalues.
The corresponding eigenvectors again match the structure given in Eq. 8, now with $n=2$, thereby validating the recursion result up to this order.
Step 3: We now have sufficient information to establish the general recursion relation $(w^{(n)},A^{(n)})\mapsto(w^{(n+1)},A^{(n+1)})$, given by,
$w^{(n+1)}_{j}$ $=\sum_{k,i}w_{k}^{(n)}v_{i}\left(A^{(n)}_{k}\right)_{i,j}^{*},$ (10)
$A^{(n+1)}$ $=\text{eig}\int\mathrm{d}h\,P(h)|w^{(n+1)}\otimes v\rangle\hskip-1.42262pt\langle w^{(n+1)}\otimes v|.$ (11)
The eigenvalue decomposition is depicted diagrammatically as,
As in Step 2, the eigenvalue decomposition of the $12\times 12$ matrix is carried out perturbatively up to $O(\tau^{4})$. The resulting eigenvectors allow us to propagate the recursion for $w^{(n)}$.
We assume the general structure of $w^{(n)}$ takes the form
$w^{(n)}=\begin{bmatrix}w_{10}^{(n)}+w_{12}^{(n)}\tau^{2}+w_{14}^{(n)}\tau^{4}\\ \mathrm{i}w_{21}^{(n)}\tau+\mathrm{i}w_{23}^{(n)}\tau^{3}\\ w_{32}^{(n)}\tau^{2}+w_{34}^{(n)}\tau^{4}\\ \end{bmatrix}$
which reflects the parity structure inherited from the expansion in $\tau$.
From the perturbative eigenvalue decomposition, we derive recursive relations for the expanding coefficients, such as,
$w_{10}^{(n+1)}=2w_{10}^{(n)},\quad w_{12}^{(n+1)}=2w_{12}^{(n)}-\frac{h^{2}}{2}w_{11}^{(n)}.$
and so on. While the full list of recursive expressions is algebraically tedious, the coefficients admit closed-form expressions:
$\begin{cases}w_{11}^{(n)}=2^{n}\\ w_{12}^{(n)}=-n2^{n}h^{2}/4\\ w_{14}^{(n)}=n2^{n}h^{2}\big{(}h^{2}(3n-1)+2\alpha^{2}(2n-1)\big{)}/96\\ w_{21}^{(n)}=-2^{n}h\sqrt{n/2}\\ w_{23}^{(n)}=2^{n}h^{3}\sqrt{n/2}(3n-1)/12\\ w_{32}^{(n)}=2^{n}\sqrt{n(2n-1)}(\alpha^{2}-3h^{2})/12\\ w_{34}^{(n)}=2^{n}\sqrt{n(2n-1)}\big{(}15h^{4}(3n-2)-15h^{2}n\alpha^{2}-2(4n-3)\alpha^{4}\big{)}/720\end{cases}$ (12)
With these renormalized vectors in hand, the matrix elements in Eq. 8 can be constructed directly from Eq. 11, completing the recursive formulation.
The key result extracted from the above procedure is the expression for the largest three eigenvalues of the reduced density matrix:
$\frac{E^{(n+1)}}{4^{n+1}}=\Big{\{}1-\frac{2(n+1)}{3}\alpha^{2}\tau^{2}+\frac{38n^{2}+43n+20}{45}\alpha^{4}\tau^{4},\frac{2(n+1)}{3}\alpha^{2}\tau^{2}-\frac{4(n+1)(3n+2)}{15}\alpha^{4}\tau^{4},\frac{4(n+1)(2n+1)}{45}\alpha^{4}\tau^{4}\Big{\}}.$ (13)
This result can be directly validated by examining the residual norm
$\|A^{(n+1)\dagger}\bar{\rho}A^{(n+1)}-E^{(n+1)}\|\sim O(\tau^{5})$
where $\bar{\rho}$ denotes the integrated density matrix appearing on the right-hand side of Eq. 11.
Moreover, Eq. 13 enables an important scaling analysis in the limit $\tau\to 0$ and $n\to\infty$ with the product $t=n\tau$ held fixed. In this limit, only the leading eigenvalue survives, growing asymptotically as $4^{n+1}$, while the subleading eigenvalues are suppressed. The relative magnitudes of the second and third eigenvalues behave as
$\frac{E^{(n+1)}_{2}}{E^{(n+1)}_{1}}\sim\frac{2}{3}\alpha^{2}\tau t,\quad\frac{E^{(n+1)}_{3}}{E^{(n+1)}_{1}}\sim\frac{8}{45}\alpha^{4}\tau^{2}t^{2},$
which is exact the results shown in the main text.
IV.2 Numerical analysis
In the previous section, by solving the recursion relation, we showed that relative values of the three largest singular values of the MPO approximation take the form $2(\alpha^{2}\tau t)/3$ and $8(\alpha^{2}\tau t)^{2}/25$. This suggests that the full singular value spectrum can be expressed as $c_{i}(\alpha^{2}\tau t)^{i},\quad i=1,2,\cdots$ where $c_{i}$ are constants.
To verify this conjecture, we numerically investigate the singular value spectrum by systematically varying the disorder strength $\alpha$, the evolution time $t$, and the Trotter step $\tau$. The results are shown as blue dots in Fig.5, with the corresponding fits to the analytic form $c_{i}(\alpha^{2}\tau t)^{i}$ indicated by red dashed lines. The data presented in Fig. 2(a) of the main text is selected from this set of results.
Overall, the numerical data agree well with the predicted form $c_{i}(\alpha^{2}\tau t)^{i}$, supporting our conjecture. To highlight the exponential decay of the singular values, we also plot the full spectrum at different times in Fig. 5(a), with each color representing a different time. At large values of $\alpha$ and $\tau$, deviations begin to appear, which can be attributed to the breakdown of the perturbative expansion, valid only in the regime $\alpha\tau\ll 1$. Nevertheless, even in this regime, the numerical singular values remain smaller than the fitted values, indicating that the form $c_{i}(\alpha^{2}\tau t)^{i}$ serves as an upper bound for the spectrum.
This combined perturbative and numerical analysis confirms that the singular values of the approximated MPO decay faster than exponentially. As a result, SeTN with a moderate bond dimension is sufficient to represent the disorder-averaged operator $\mathrm{O}$ reliably, even at long times.
This scaling behavior also has a clear physical interpretation: as the number of layers $n$ increases, the size of the effective Hilbert space grows exponentially, but the relevant physical information remains confined to a low-entanglement subspace that encodes the same continuous-time dynamics. The dimensionless parameter $\alpha^{2}t^{2}$ characterizes the total information content from disorder averaging, while the information encoded per layer scales as $\alpha^{2}t^{2}/n=\alpha^{2}\tau t$, which directly controls the singular value spectrum.
From a numerical perspective, this analysis reveals a key trade-off: decreasing the time step $\tau$ improves the accuracy of the approximation and reduces the required bond dimension, but also increases the Trotter number $n=t/\tau$, which lengthens the MPO and raises computational cost. Therefore, maintaining the condition $\alpha^{2}\tau t\ll 1$ provides a practical guideline for choosing $\tau$. In particular, to ensure efficient SETN compression, one should use $n\gg\alpha^{2}t^{2}$ or equivalently $\tau\ll 1/(\alpha^{2}t)$.
[…]
IV Spectral form factor of time-independent Hamiltonian
In this section, we clarify the reasoning behind our choice of the spectral form factor (SFF) definition used in the main text, especially for autonomous (time-independent) systems. Although many studies do not explicitly address the subtleties of SFF definitions, a detailed discussion helps clarify the necessity of unfolding and filtering procedures, and the limitations of approximations often used.
For Floquet systems, where the dynamics is governed by a unitary Floquet operator $F$ with eigenphases $\{\phi_{i}\}$, the spectral density is
$\varrho(\phi)=\frac{1}{D}\sum_{i=1}^{D}\delta(\phi-\phi_{i})=\frac{1}{2\pi D}\sum_{n=-\infty}^{\infty}t_{n}\mathrm{e}^{\mathrm{i}n\phi},$
where,
$t_{n}=\sum_{i=1}^{D}\mathrm{e}^{\mathrm{i}n\phi_{i}}=\operatorname{tr}F^{n}.$
The connected two-point correlation function is defined as
$S(\phi,\phi^{\prime})\equiv\overline{\varrho(\phi)\varrho(\phi^{\prime})}-\overline{\varrho(\phi)}\;\overline{\varrho(\phi^{\prime})},$
where the overline $\overline{O}=\int\mathrm{d}^{D}\{\phi_{i}\}\,P(\{\phi_{i}\})O$ ensemble or disorder averaging. Assuming exchange symmetry among the eigenphases which is true in Floquet system, it follows that
$\overline{t_{n}t_{n^{\prime}}}=\delta_{n^{\prime},-n}\overline{|t_{n}|^{2}},\quad\quad\overline{\varrho(\phi)}=1/2\pi.$
Then, the two-point correlation function becomes,
$S(\phi,\phi^{\prime})=\frac{1}{2\pi^{2}D^{2}}\sum_{n=1}^{\infty}\overline{|t_{n}|^{2}}\cos\left(n(\phi-\phi^{\prime})\right).$
and the spectral form factor is defined as the Fourier coefficients:
$K(n)=\overline{|t_{n}|^{2}}=\overline{\sum_{i,j=1}^{D}\mathrm{e}^{\mathrm{i}n(\phi_{i}-\phi_{j})}}.$
For time-independent Hamiltonians $H$, the eigenphases $\{\phi_{i}\}$ are replaced by eigenenergies $\{E_{i}\}$, which lack exchange symmetry and generally exhibit a non-uniform local spectral density. In practice, to extract meaningful correlations (and thus determine timescales like the Thouless time), it is common practice to unfold the spectrum, rescaling eigenenergies $E_{i}\to\epsilon_{i}$ to achieve a uniform average spacing [gomezMisleadingSignaturesQuantum2002, haakeQuantumSignaturesChaos2018, gharibyanOnsetRandomMatrix2018a, suntajsQuantumChaosChallenges2020b, sierantPolynomiallyFilteredExact2020]. Additionally, edge effects due to finite bandwidth are mitigated by applying a smooth filter function $\rho(\epsilon)$ (e.g., a Gaussian):
$K(t)=\overline{\left|\sum_{i=1}^{D}\rho(\epsilon_{i})\mathrm{e}^{-2\pi\mathrm{i}t\epsilon_{i}}\right|^{2}},$ (14)
where $\{\epsilon_{1}\leq\epsilon_{2}\leq\cdots\leq\epsilon_{D}\}$ are the unfolded eigenvalues.
This definition of the SFF is essential for accurately determining the Thouless time in autonomous systems, as it accounts for the non-uniform local density of states and mitigates edge effects. However, as argued in [gharibyanOnsetRandomMatrix2018a], sample-to-sample fluctuations in the eigenvalue density do not qualitatively alter the ramp structure of the SFF. The primary impact is a rounding of the sharp ramp-to-plateau transition, as different spectral regions saturate at different times. The magnitude of this effect depends on the spectral density $\varrho(E)$ and the specific system studied.
This robustness of the ramp implies that density fluctuations across samples do not significantly affect the overall ramp structure, making the SFF a reliable diagnostic of the Thouless time even without unfolding. For simplicity, we adopt the SFF definition given in the main text and defer a detailed analysis of unfolding and filtering procedures to future work.
V Chaotic signature in disordered transverse-field Ising model
VI.1 Mean Level Spacing Ratio
An important feature of chaotic systems is the presence of spectral correlations, which can be quantified by the level spacing ratio $\langle r\rangle$ [atasDistributionRatioConsecutive2013a]:
$r=\min\left\{r_{n},\frac{1}{r_{n}}\right\},\quad r_{n}=\frac{E_{n+1}-E_{n}}{E_{n}-E_{n-1}},$ (15)
where $\{E_{n}\}$ denotes the ordered eigenvalues of $H_{\text{dTFIM}}$. The average value $\langle r\rangle$, computed over all eigenenergies, takes a characteristic value $r_{\text{GOE}}\simeq 0.5307$ for systems with time-reversal symmetry exhibiting Gaussian orthogonal ensemble (GOE) statistics, and $r_{\text{PS}}\simeq 0.3863$ for systems following Poisson statistics, typical of integrable or localized behavior.
The level spacing ratio $\langle r\rangle$ is shown in the main panel of Fig. 6 (a) as a function of disorder strength $\alpha$ for system size $L=11,13$ and $15$. It remains close to the GOE value $r_{\text{GOE}}$ in the range $0.2\sim 1.3$, indicating chaotic behavior, and approaches the Poisson value $r_{\text{PS}}$ for strong disorder, $\alpha\gtrsim 2.5$, signaling localization. Here and in main text, we focus on the representative case of $\alpha=0.5$, marked by the red star in the main panel.
For small disorder $\alpha<0.1$ as shown in the inset of Fig. 6 (a), the level spacing ratio deviates from both limiting values. This can be attributed to approaching the clean transverse field Ising model, where hidden symmetries need to be resolved before applying level statistics reliably. We also observe that the chaotic region slightly broadens with increasing system size, consistent with previous findings in other time-independent Hamiltonian systems [morningstarAvalanchesManybodyResonances2022a]. This trend suggests that quantum chaos remains robust in the central part of the disorder range.
VI.2 Thouless Time Scaling
Within RMT, the GOE SFF exhibits a universal ramp-plateau structure [liuSpectralFormFactors2018b]:
$K_{\text{GOE}}(t)=\begin{cases}2t-t\log(1+2t/D),&t\leq D\\ 2D-t\log\frac{2t/D+1}{2t/D-1},&t>D\end{cases},$ (16)
where $D$ is the Hilbert space dimension, and $t_{\text{H}}=D$ denotes the Heisenberg time.
The ramp-plateau structure signals spectral correlations indicative of quantum chaos. The ramp’s onset time $t_{\text{Th}}$, scaling with system size, reflects an intrinsic dynamical timescale characterized by Lyapunov exponents [bertiniExactSpectralForm2018, bertiniRandomMatrixSpectral2021a, chanSpectralLyapunovExponents2021].
For systems without conserved quantities, the growth of $t_{\text{Th}}$ arises from domain walls in emergent statistical mechanical models, which separate expanding RMT-like regions [chanSpectralStatisticsSpatially2018c, garrattLocalPairingFeynman2021c, garrattManyBodyDelocalizationSymmetry2021a]. In contrast, for systems with conserved quantities or local constraints, the scaling behavior of $t_{\text{Th}}$ is governed by the diffusive or subdiffusive transport of the conserved charges [friedmanSpectralStatisticsManyBody2019a, moudgalyaSpectralStatisticsConstrained2021a, dagManybodyQuantumChaos2023a].
Figure 6: (Color online) Spectral form factor (SFF) of the disordered transverse field Ising model for various system sizes, zoomed into the region near the Thouless time. Each curve is averaged over 2440 disorder realizations of $\boldsymbol{h}$, uniformly distributed with disorder strength $\alpha=0.5$ and coupling constants $J=b=1$. The shaded region represents a linear fit of the data between $t=20$ and $t=90$, with a 95% confidence interval. The inset (b) shows the extracted Thouless time (indicated by the black circle in the main panel) as a function of system size.
The zoomed-in SFF of dTFIM with different lengths is plotted in Fig. 6 (b), where we maximally calculate 2440 averages of realizations in the systems with Hilbert space dimension $2^{16}=65536$.
We fit the SFF in the region between $20$ and $90$ with a linear function. $95\%$ confidence interval of the fit is shown as the shaded region. We use the first section with the upper bound of the shaded region after the peak as an estimation of the Thouless time. Even though it is not the most accurate way to determine the Thouless time, it is sufficient to show the log scale increase as the system size, with is consistent with many other models without conserved quantities. The behavior of $K(t)$ at intermediate times $1\ll t\lesssim t_{\text{Th}}$ also remains an important open problem in quantum chaos.
VI Implementation of Transfer Matrix Methods
To compute the leading eigenvalue of the SeTN-derived transfer matrix, we employ both the Arnoldi and non-Hermitian DMRG methods. In this section, we outline the technical details of our implementation, with an emphasis on exploiting the specific tensor network structure of the disordered transverse-field Ising model (disordered TFIM) after Trotter decomposition.
VII.1 Efficient Application of the Transfer Matrix
The evolution operator of the dTFIM decomposes into a structure identical to that of the kicked Ising model, allowing for efficient tensor contractions. The clean part of the Hamiltonian, $H_{\text{clean}}$, comprises an $x$-field term and a nearest-neighbor $zz$ interaction term. The former yields on-site gates, while the latter is a diagonal two-site gate that can be simplified to an effective on-site gate in the time direction:
$(\mathrm{e}^{-\mathrm{i}\tau J\sigma^{z}_{i}\sigma^{z}_{i+1}})_{j_{n}^{[i]}j_{n+1}^{[i]},j_{n}^{[i+1]}j_{n+1}^{[i+1]}}$ $=\mathrm{e}^{-\mathrm{i}\tau J\sigma_{j^{[i]}_{n}}\sigma_{j^{[i+1]}_{n}}}\delta_{j_{n}^{[i]}j_{n+1}^{[i]}}\delta_{j_{n}^{[i+1]}j_{n+1}^{[i+1]}}$
$=\sum_{k_{n}^{[i]},k_{n}^{[i+1]}}\delta_{j_{n}^{[i]}j_{n+1}^{[i]}k^{[i]}_{n}}\mathrm{e}^{-\mathrm{i}\tau J\sigma_{k^{[i]}_{n}}\sigma_{k^{[i+1]}_{n}}}\delta_{j_{n}^{[i+1]}j_{n+1}^{[i+1]}k^{[i+1]}_{n}}$,
where $\sigma_{j^{[i]}_{n}}=1$ if $j^{[i]}_{n}=1$ and 1-1 if $j^{[i]}_{n}=2$. This structure is illustrated in Fig. 7(a), with $(u_{1})_{k^{[i]}_{n}k^{[i+1]}_{n}}=\mathrm{e}^{-\mathrm{i}\tau J\sigma_{k^{[i]}_{n}}\sigma_{k^{[i+1]}_{n}}}$. The resulting $W$ tensor can thus be factorized into a five-index Kronecker delta, a time-direction gate $u_{1}$, and a space-direction gate $u_{2}$, as shown in Fig. 7(b).
Figure 7: (Color online) (a) Decomposition of the zzzz interaction gate into Kronecker deltas and a time-direction gate $u_{1}$. (b) Simplified form of the $W$ tensor as a combination of local gates and delta tensors.
Using this structure, we represent the transfer matrix as in Fig. 8(a), enabling efficient application to a vector without explicitly forming the full transfer matrix. The matrix-vector multiplication proceeds in stages: first applying $u_{1}$ and $u_{1}^{*}$, followed by alternating contractions with the disorder-encoding tensor $u_{3}$ and the evolution tensors $u_{2},u_{2}^{*}$. This process is visualized in Fig. 8(c). In the standard tensor network compression implementation via QR-SVD, we drop the component corresponding to the singular values with relative magnitude between $10^{-14}$. The locality of the gates allows for efficient implementation of the Krylov-Schur method to extract the dominant eigenvalues.
Figure 8: (Color online) (a) Structure of the transfer matrix contraction. (b) MPO representation of the variational state in DMRG. (c) Efficient contraction procedure for transfer matrix application. (d) Local effective Hamiltonian for the DMRG update.
VII.2 Non-Hermitian DMRG with Eigenvalue Tracking
For the non-Hermitian DMRG approach, we similarly exploit the MPO structure of the transfer matrix. Representing the variational state as an MPO [Fig. 8(b)], the local effective Hamiltonian appears as in Fig. 8(d). As with the Arnoldi method, the contraction proceeds gate by gate, leveraging locality for efficiency.
Standard DMRG typically uses the Lanczos method to diagonalize the effective Hamiltonian. However, since the transfer matrix is non-Hermitian (due to $\tau\ll 1$ and deviation from the self-dual point), we instead use the Arnoldi method for each update. At longer times, we observe the appearance of eigenvalues with similar magnitude but differing complex phases, which can severely impact DMRG convergence if selected randomly. To address this, we implement an eigenvalue tracking scheme: at each DMRG step, we select the eigenvalue closest (in Euclidean distance) to the one chosen in the previous iteration, provided its magnitude is also larger. This continuity improves convergence significantly.
With this optimization, we successfully compute up to $t=1.8$ (with $n=40$ Trotter steps), shown in Fig. 3(b) in main text, corresponding to a state space dimension of $2^{80}$. For longer times, convergence deteriorates, likely due to near-degeneracies or the proliferation of quasi-degenerate eigenvalues with different phases. Further study is needed to fully resolve these issues.
VII A toy model for Thouless time scaling and RMT ramp
As discussed in the main text, we conjecture that around the Thouless time, the leading eigenvalues of the effective transfer matrix gradually become nearly degenerate over time, generating the linear ramp in the spectral form factor (SFF). At fixed time, the true largest eigenvalue eventually dominates as the system size increases. To illustrate this idea, we introduce a simple toy model that captures the essential features of both the emergence of the RMT ramp and the growth of the effective Thouless time.
We consider a set of eigenvalues $\{\lambda_{i}(t)\}$ of an effective transfer matrix at time $t$,
$\lambda_{i}(t)=\begin{cases}1,&\text{for}\;i=0\\ 1-0.99^{\,t-i},&\text{for}\;0<i\leq t\\ 0,&\text{for}\;i>t\end{cases}$ (17)
where the subleading eigenvalues gradually approach the leading one, mimicking the flattening of the spectrum around the Thouless time, while eigenvalues with $i>t$ are set to zero to limit the number of contributing modes. From these eigenvalues, we define the corresponding SFF for system size $L$ as
$K(t)=\sum_{i}\lambda_{i}(t)^{L}.$ (18)
Figure 9 illustrates the behavior of this toy model. Panel (a) shows the spectrum $\lambda_{i}(t)$ as a function of $t$, highlighting the growing near-degeneracy of the leading eigenvalues. Panel (b) shows $K(t)$ for system sizes $L=9$ to $20$, demonstrating two key features: (i) a linear ramp in $t$, reminiscent of the RMT prediction, and (ii) the growth of the transient time at which the single leading eigenvalue ceases to dominate (black dashed line), reminiscent of the Thouless time behavior shown in Fig. 6(b). Despite its simplicity, this toy model captures the qualitative features of the transition from the regime dominated by a single eigenvalue to the regime where the RMT-predicted ramp emerges, illustrating our conjecture in a clear and intuitive way.
Figure 9: Toy model illustrating the conjectured transition. (a) Eigenvalues $\lambda_{i}(t)$ as a function of $t$, with $\lambda_{0}=1$ and subleading eigenvalues $\lambda_{i}(t)=1-0.99^{t-i}$ for $0<i\leq t$. (b) Corresponding spectral form factor $K(t)=\sum_{i}\lambda_{i}^{L}$ for system sizes $L=9$ to 2020 (top to bottom), showing a linear ramp in $t$ and the growth of the effective Thouless time (black dashed line), defined as the time when subleading eigenvalues start to contribute significantly. This toy model is intended to illustrate qualitative features rather than provide quantitative predictions.