### v2

# Statistics-encoded tensor network approach in disordered quantum many-body spin chains

Abstract
Simulating the dynamics of quantum many-body systems with disorder is a fundamental challenge. In this work, we propose a general approach—the statistics-encoded tensor network (SeTN)—to study such systems. By encoding disorder into an auxiliary layer and averaging separately, SeTN restores translational invariance, enabling a well-defined transfer matrix formulation. We derive a universal criterion, $n\gg\alpha^{2}t^{2}$, linking discretization $n$, disorder strength $\alpha$, and evolution duration $t$. This sets the resolution required for faithful disorder averaging and shows that encoding is most efficient in the weak-disorder, typically chaotic regime. Applied to the disordered transverse-field Ising model, SeTN shows that, over the numerically accessible time window, the spectral form factor is governed by the leading transfer-matrix eigenvalue, in contrast to the kicked Ising model. SeTN thus provides a novel framework for probing the disorder-driven dynamical phenomena in many-body quantum systems.
I Introduction
Simulating quantum many-body systems with disorder is crucial for understanding a wide range of phenomena, including quantum chaos [gutzwillerChaosClassicalQuantum1990a, dalessioQuantumChaosEigenstate2016b, haakeQuantumSignaturesChaos2018], many-body localization (MBL) [nandkishoreManyBodyLocalizationThermalization2015c, abaninColloquiumManybodyLocalization2019b, sierantManybodyLocalizationAge2025], quantum spin glass [brayReplicaTheoryQuantum1980, sachdevGaplessSpinfluidGround1993], and criticality [fisherCriticalBehaviorRandom1995, pichCriticalBehaviorGriffithsMcCoy1998]. Recently, the chaotic dynamics of disordered systems predicted by random matrix theory (RMT) has attracted growing interest, driven by advances in experimental techniques [liEmergentUniversalQuench2024, dongMeasuringSpectralForm2025].
Analytical progress on disorder-averaged dynamics has primarily been achieved in random quantum circuits and dual-unitary models, where Haar averaging or self-duality enables exact transfer-matrix treatments [chanSolutionMinimalModel2018c, bertiniExactSpectralForm2018, vonkeyserlingkOperatorHydrodynamicsOTOCs2018a, zhouEmergentStatisticalMechanics2019, nahumOperatorSpreadingRandom2018c, sonnerCharacterizingManybodyLocalization2022, riddellStructuralStabilityHypothesis2024, bertiniExactlySolvableManybody2025a]. However, these approaches cannot be directly extended to generic time-independent Hamiltonians, where self-duality is absent and the transfer-matrix dimension grows exponentially, hindering practical implementation. Numerically, exact diagonalization (ED) is restricted to small system sizes [pietracaprinaShiftinvertDiagonalizationLarge2018a, sierantPolynomiallyFilteredExact2020a], while conventional tensor-network methods and strong-disorder renormalization approaches [doggenManybodyLocalizationDelocalization2018a, chandaTimeDynamicsMatrix2020, hikiharaNumericalRenormalizationgroupStudy1999, goldsboroughSelfassemblingTensorNetworks2014a] treat each disorder realization independently and are applicable mostly to the deep MBL regime. Another widely used route in this regime, quantum-parallel method [alvarezQuantumParallelismTool2008a, paredesExploitingQuantumParallelism2005c, andraschkoPurificationManyBodyLocalization2014, hubigTimedependentStudyDisordered2019a], embeds disorder realizations taking only discrete values into an auxiliary system and evolves them in parallel. However, when the disorder follows a continuous distribution, as is typical in studies of chaotic dynamics, the method requires an auxiliary space of infinite dimension.
In this letter, we propose the statistics-encoded tensor network (SeTN), a general approach for studying disordered many-body systems with time-independent Hamiltonians. In SeTN, we encode disorder into an auxiliary layer on which the average is taken independently at each site (Fig. 1(b)). By analyzing the decay rate of the singular values, we derive a universal criterion for the temporal resolution required for efficient disorder encoding: $n\gg\alpha^{2}t^{2}$, where $n$ is the number of time steps, $\alpha$ the disorder strength, and $t$ the evolution duration. This criterion shows that continuous disorder distributions can be efficiently encoded, particularly in the weak-disorder regime, making disorder-averaged transfer matrices numerically feasible for generic Hamiltonians.
After averaging, SeTN restores spatial translation invariance, yielding a compact MPO transfer matrix that can be directly analyzed using Krylov iteration [stewartKrylovSchurAlgorithmLarge2002] and density matrix renormalization group (DMRG) methods [whiteDensityMatrixFormulation1992a, whiteDensitymatrixAlgorithmsQuantum1993]. As a demonstration, we investigate the spectral form factor (SFF) of the disordered transverse-field Ising model in its chaotic regime, revealing, within the numerically accessible regime, a prolonged pre-RMT transient dominated by a single, non-degenerate leading eigenvalue of the transfer matrix. This pathway to RMT behavior contrasts with Floquet models such as the kicked Ising model, where chaotic spectral statistics and eigenvalue degeneracies appear from the outset [chanSolutionMinimalModel2018c].
II Statistics-encoded tensor network approach
We consider a quantum many-body system with local disorder, described by the Hamiltonian $\mathcal{H}[\boldsymbol{h}]=H+\sum_{i=1}^{L}h_{i}H_{i}$, where $H$ is the disorder-free part, $\{H_{i}\}$ are local operators acting on site $i$, $\{h_{i}\}$ are disorder variables which are independent and identically distributed, and $L$ is the system size. Throughout this work we set $\hbar=1$. To simulate the real-time dynamics, we apply a first-order Trotter decomposition with a small time step $\tau\ll 1$, which gives the evolution operator $U(\tau)=\mathrm{e}^{-\mathrm{i}\tau H}\mathrm{e}^{-\mathrm{i}\tau\sum_{i}h_{i}H_{i}}+O(\tau^{2})$. This decomposition enables a TN representation of the full evolution operator (see Fig. 1(a)). For simplicity, we focus on the widely studied case where $H=\sum_{i}H_{i,i+1}$ includes only nearest-neighbor interactions and $H_{i}=\sigma^{z}_{i}$ is the Pauli-$z$ operator on site $i$. Generalizations to longer-range interactions or different disorder couplings are straightforward [supplementaryMaterial].
To proceed, as shown in Fig. 1(a), we first decompose the two-body unitary gates into two local tensors,
$\langle j_{n}^{[i]}j_{n+1}^{[i]}|\mathrm{e}^{-\mathrm{i}\tau H_{i,i+1}}|$ $j_{n}^{[i+1]}j_{n+1}^{[i+1]}\rangle=V^{(L)}_{j_{n}^{[i]}j_{n+1}^{[i]},k_{n}^{[i]}}V^{(R)}_{k_{n}^{[i]},j_{n}^{[i+1]}j_{n+1}^{[i+1]}},$ (1)
where $j_{n}^{[i]}$ refers to site $i$ and time step $n$, with repeated indices summed over implicitly. This can be achieved via singular value decomposition (SVD) or spectral decomposition. At the same time, we factor out the diagonal disorder term using a delta tensor: $\langle j_{n}^{[i]}|\mathrm{e}^{-\mathrm{i}\tau\sigma^{z}_{i}}|j_{n+1}^{[i]}\rangle=v_{l_{n}^{[i]}}\delta_{l_{n}^{[i]}j_{n}^{[i]}j_{n+1}^{[i]}}$, where $\delta_{ijk}$ is the three-order Kronecker tensor (equal to $1$ when $i=j=k$ and 0 otherwise), and $v_{l}=\mathrm{e}^{-\mathrm{i}\tau h_{i}l}$ with $l\in\{1,-1\}$ encodes the effect of local disorder. Throughout this work, we use the indices $j,k$ and $l$ to denote the time, space, and layer directions, respectively, as indicated by the coordinate system in Fig. 1. By combining these decompositions, we define a local tensor $W$ as
$W_{j_{n}^{[i]}j_{n+1}^{[i]},k_{n}^{[i-1]}k_{n}^{[i]},l_{n}^{[i]}}=\delta_{l_{n}^{[i]}j_{n}^{[i]}p_{n}^{[i]}}V^{(R)}_{k_{n}^{[i-1]},p_{n}^{[i]}q_{n}^{[i]}}V^{(L)}_{q_{n}^{[i]}j_{n+1}^{[i]},k_{n}^{[i]}}$, (2)
where $p$ and $q$ are auxiliary indices used for internal contractions. This construction is illustrated in the third column of Fig. 1(a), where the spatial slice at site $i$ corresponds to $\prod_{n}v_{l_{n}^{[i]}}W_{j_{n}^{[i]}j_{n+1}^{[i]},k_{n}^{[i-1]}k_{n}^{[i]},l_{n}^{[i]}}$.
Figure 1: (Color online) (a) Tensor network (TN) representation of the Trotterized evolvation operator. The tensor encoding the disorder is decomposed into a vector (blue circle) and a Kronecker delta tensor (black point). (b) A general two-layer TN representation of the disorder-averaged transfer matrix $\mathrm{T}$. The orange-shaded region indicates the disorder average. (c) Diagrammatic illustration of the SeTN compression. $S^{(n)}$ is the singular values obtained via corresponding singular value decomposition.
Figure 2: (Color online) (a) Blue dots: squared relative singular values from SeTN compression under various parameters (Num.). Dashed lines: analytic predictions from perturbation theory (Ana.), with coefficients fitted to numerical data. (b) Blue dots: first seven coefficients $c_{i}$ obtained by fitting the data in (a). These are further fitted to exponential (green dashed), Gaussian (purple dashed), and gamma-like (red dot-dashed) functions. (c) Spectral form factor of the disordered TFIM with system size $L=4$, computed using four methods: exact diagonalizing 1000 disorder realizations (avg #1000, green dotted; shaded band indicates the standard error), numerical integration (Exact, blue dashed), SeTN (red solid), and perturbative results (Pert., gray dot-dashed). The mean deviation between SeTN and the exact result is approximately 0.021. Parameters: $J=b=1$, Trotter step $\tau=0.005$, averaging over $M=10^{6}$ realizations; singular values below $10^{-10}$ are truncated.
We study disorder-averaged dynamics, for which many quantities of interest including averaged operator expectation values $\langle\!\langle O(t)\rangle\!\rangle=\langle\,\operatorname{Tr}[O\,U(t)\rho(0)U^{\dagger}(t)]\,\rangle$, as well as the SFF $K(t)=\langle\,\operatorname{Tr}[U(t)\otimes U^{\dagger}(t)]\,\rangle$, naturally admit a two-layer tensor-network representation obtained by contracting the forward and backward evolutions $U$ and $U^{\dagger}$ (dark/light blue in Fig. 1(b)). The effect of disorder is encoded in a statistics layer (orange region) which, after averaging, renders all spatial sites equivalent. This structure motivates the definition of a transfer matrix $\mathrm{T}$ that captures the combined action of unitary dynamics and disorder averaging across a single spatial slice,
$\mathrm{T}_{k_{1}^{\prime[i-1]},...,k_{n}^{\prime[i-1]},k_{1}^{[i-1]},...,k_{n}^{[i-1]};k_{1}^{\prime[i]},...,k_{n}^{\prime[i]},k_{1}^{[i]},...,k_{n}^{[i]}}$ (3)
$=\Big\langle\prod_{p=1}^{n}W^{*}_{j_{p}^{\prime[i]}j_{p+1}^{\prime[i]},k_{p}^{\prime[i-1]}k_{p}^{\prime[i]},l_{p}^{\prime[i]}}v^{*}_{l_{p}^{\prime[i]}}v_{l_{p}^{[i]}}W_{j_{p}^{[i]}j_{p+1}^{[i]},k_{p}^{[i-1]}k_{p}^{[i]},l_{p}^{[i]}}\Big\rangle$
$=\mathrm{O}_{l^{\prime}_{1},...,l^{\prime}_{n};l_{1},...,l_{n}}\prod W^{*}W,$
where the indices on the conjugated tensors are distinguished by primes and $\langle{\cdot}\rangle$ denotes the disorder average.
In the last line, We separate the disorder-averaged tensor $\mathrm{O}$ from the unitary layers (omitting site superscripts and tensor subscripts for clarity), which can be obtained explicitly by integrating over the distribution of $h$:
$\mathrm{O}_{l^{\prime}_{1},...,l^{\prime}_{n};l_{1},...,l_{n}}\equiv\Big\langle{\prod_{p=1}^{n}v_{l^{\prime}_{p}}^{*}v_{l_{p}}}\Big\rangle=\int\mathrm{d}h\;\mathrm{e}^{\mathrm{i}h\tau\sum_{p}(l^{\prime}_{p}-l_{p})}P(h)$ (4)
[…]
where $P(h)$ denotes the normalized probability density function of the disorder variable $h$, i.e., $\int dh\,P(h)=1$. We give explicit formulas for both uniform $\mathcal{U}(-\alpha,\alpha)$ and Gaussian $\mathcal{N}(0,\sigma^{2})$ disorder. While a similar expression for the Gaussian case was derived in the KIM context [bertiniExactSpectralForm2018], the present formulation applies to general static Hamiltonians with arbitrary on-site disorder. Moreover, when $P(h)$ is discrete, $\mathrm{O}$ reduces exactly to the standard quantum-parallel encoding [supplementaryMaterial].
Although Eq. 4 represents the ideal analytic average, directly constructing the $2^{n}\times 2^{n}$ tensor $\mathrm{O}$ is numerically infeasible. In practice, we therefore adopt a finite-sample approximation with $M$ disorder realizations sampled from the distribution $P(h)$, and compress the resulting MPO through the two-step procedure illustrated in Fig. 1(c). First, each Se-Layer realization is written as a matrix product operator (MPO) with bond dimension 1; summing over $M$ such realizations via MPO addition [crosswhiteFiniteAutomataCaching2008] yields an MPO of bond dimension $M$ encoding the disorder average. Second, we iteratively compress this MPO using SVD. At $n=1$ (pink box in Fig. 1(c)), the singular values $S^{(1)}$ below a fixed threshold are discarded. The truncated tensors are contracted with $A^{(2)}$ and decomposed again (blue box in Fig. 1(c)), and the process continues site by site to the final layer.
Notely, since each step involves only local operations and the tensors $A^{(n)}$ are identically constructed via MPO addition from individual realizations, the overall memory requirement reduces to $O(M)$ [supplementaryMaterial], compared to the naive $O(nM^{2})$ scaling. This significantly enhances computational efficiency and enables averaging over a large number of realizations ($M\gg 1$) in practice.
III Properties of Se-Layer and SeTN effectiveness
The previous section outlined the implementation of SeTN. A natural question is whether disorder averaging can be faithfully encoded into the TN representation. This effectiveness crucially depends on the behavior of the singular value spectrum $S^{(n)}$, illustrated in Fig. 1(b): only when the singular values decay rapidly, ideally exponentially, can the SeTN accurately compress the averaged transfer matrix with a low bond dimension. To analyze this, we derive an analytic expression for the spectrum by solving the recursive structure in SeTN compression perturbatively [supplementaryMaterial]:
$\left(S_{i}(t)/S_{0}(t)\right)^{2}\simeq c_{i}\left(\alpha^{2}\tau t\right)^{i},$ (5)
where $S_{i}(t)$ denotes the $i$-th singular value at the $n$-th compression step (with $n=t/\tau$), ordered in descending magnitude, $S_{0}(t)$ is the largest and $\alpha$ is the disorder strength. The coefficients $c_{i}$ depend only on the disorder statistics; for a uniform distribution, the first two values can be computed exactly as $c_{1}=2/3$ and $c_{2}=8/45$ [supplementaryMaterial]. This result shows that the decay of the singular values is governed by the dimensionless parameter $\alpha^{2}\tau t$, which confirms the exponential suppression of subleading components in the singular value spectrum when $\alpha^{2}\tau t\ll 1$, or equivalently,
$n\gg\alpha^{2}t^{2}.$ (6)
This condition ensures that SeTN achieves accurate, low-bond-dimension compression of disorder-averaged transfer matrices.
To validate this analytic result, we numerically investigate the singular value spectra under various parameter settings [supplementaryMaterial] and compare them with the perturbative result. As shown in Fig. 2(a), the blue dots represent the normalized squared singular values $(S_{i}(t)/S_{0}(t))^{2}$ obtained from numerical simulations using a uniform disorder distribution, while the red dashed lines show the corresponding analytic estimates. The excellent agreement between the numerical data and theoretical predictions in the region $\alpha^{2}\tau t\ll 1$ confirms the exponential decay behavior discussed above.
In Fig. 2(b), we further analyze the coefficients $c_{i}$ appearing in Eq. 6. The plot shows the first seven coefficients extracted by fitting the numerical spectra in Fig. 2(a) to the analytic form. Notably, these coefficients decay faster than exponentially but slower than Gaussian, and are well approximated by a gamma-like function, which highlights the intrinsic efficiency of the SeTN method. Importantly, despite the rapid decay, the leading coefficients remain large. For a typical setting where $\alpha^{2}\tau t=10^{-2}$ (e.g., $\alpha=0.1$, $\tau=0.01$, $t=100$), the corresponding contribution to $(S_{i}/S_{0})^{2}$ at $i=7$ is on the order of $10^{-14}$, several orders of magnitude smaller than the coefficient $c_{7}=10^{-6}$. Thus, the overall compression efficiency of SeTN is still primarily governed by the dimensionless combination $\alpha^{2}\tau t$, as emphasized earlier.
For fixed disorder strength $\alpha$ and time $t$, we provide a physical interpretation of the above properties: while increasing the discretization number $n$ leads to exponential growth in the dimension of the transfer matrix, the total information introduced by disorder averaging, quantified by the dimensionless parameter $\alpha^{2}t^{2}$, remains unchanged. In the SeTN approach, this information is distributed across a tensor network of length $n$, such that the amount of information encoded per time step scales as $\alpha^{2}t^{2}/n=\alpha^{2}\tau t$. This local information density is directly related to the decay of the singular value spectrum, which allows efficient compression and accurate representation of disorder-averaged dynamics when $n\gg\alpha^{2}t^{2}$.
Figure 3: (Color online) (a) SFF of the disordered TFIM with sizes $L=9$ to 1616 (top to bottom), computed by ED using 2440 independent disorder realizations per size. Standard errors are below 4% at all times, and are therefore not visible on the logarithmic scale. Shaded regions (i) and (ii) indicate the first and second rebounce windows. (b) Magnitudes of the leading and subleading eigenvalues of the SeTN-derived transfer matrix as functions of time, obtained using Krylov (blue crosses) and non-Hermitian DMRG (red circles). Horizontal bars (i) and (ii) match the time windows in (a), highlighting the near-degeneracy of dominant eigenvalues. Black lines represent the fitted coefficient extracted by fitting the data in panel (a) to the form $\lambda(t)^{L}$, and the shaded region indicates the 80% confidence interval. Parameters: Trotter step $\tau=0.05$, $J=b=1,\alpha=0.5$.
IV Scaling behavior of spectral form factor
The preceding analysis shows that SeTN is more efficient for weak disorder, as singular values decay faster. In this regime, the system is typically driven into chaos, which is challenging for conventional tensor networks. To illustrate SeTN’s ability to handle chaotic dynamics, we investigate the spectral form factor (SFF) of the disordered transverse-field Ising model (disordered TFIM) in its chaotic regime.
The SFF, defined as the Fourier transform of the two-point correlation function of the spectral density [supplementaryMaterial],
$K(t)=\langle{|\operatorname{Tr}U(t)|^{2}}\rangle=\operatorname{Tr}\langle{U(t)\otimes U^{*}(t)}\rangle,$ (7)
where $U(t)=\mathrm{e}^{-\mathrm{i}Ht}$ is the evolution operator, serves as one of the simplest analytically tractable probes of quantum chaos. It captures both short- and long-range spectral correlations and has been extensively used in many-body chaos studies [chanSolutionMinimalModel2018c, kosManyBodyQuantumChaos2018b, chanSpectralStatisticsSpatially2018c, bertiniExactSpectralForm2018, dongMeasuringSpectralForm2025], where it typically exhibits an initial transient followed by the ramp–plateau structure predicted by RMT [haakeQuantumSignaturesChaos2018, liuSpectralFormFactors2018b].
Here, we study the SFF in the disordered TFIM whose Hamiltonian reads, $\mathcal{H}[\boldsymbol{h}]=J\sum_{i=1}^{L-1}\sigma^{z}_{i}\sigma^{z}_{i+1}+b\sum_{i}^{L}\sigma^{x}_{i}+\sum_{i}h_{i}\sigma^{z}_{i}$, where $J$ is the coupling strength, $b$ is the magnetic field, and $h_{i}$ are independent random variables drawn from a uniform distribution in the range $\alpha,\alpha]$. We use open boundary conditions throughout; the results are insensitive to the choice of boundary condition. Setting $J=b=1$ and $\alpha=0.5$ where the system is chaotic [supplementaryMaterial], we compute the SFF via ED for $L=9$ to $16$ averaging $2440$ independent disorder realizations for each size. The results clearly display the onset of the linear ramp as shown in Fig. 3(a).
Since the SFF is non-self-averaging [prangeSpectralFormFactor1997b], reliable evaluation typically requires extensive disorder averaging. By expressing the SFF into a two-layer tensor-network form, however, SeTN can be directly employed as outlined in the previous section. Fig. 2(c) shows that SeTN (red line) reproduces converged results obtained from numerical integration (blue dashed line) even at long times. For comparison, we include disorder-averaged ED ($10^{3}$ samples, green dotted) and second-order perturbation results (SeTN with bond dimension three [supplementaryMaterial], gray dash-dotted). The slight ED deviation (see inset of Fig. 2(c)) highlights SeTN’s accuracy, whereas the breakdown of perturbation theory emphasizes the importance of small singular values for long-time dynamics.
A key advantage of SeTN is its ability to restore spatial translational invariance broken by disorder, thereby enabling a well-defined transfer matrix and access to the thermodynamic limit. In Fig. 3(b), we show the largest eigenvalues of the transfer matrix, computed by Krylov iteration (blue crosses) and non-Hermitian DMRG (red circles) [supplementaryMaterial]. For reference, we fit the finite-size ED data in Fig. 3(a) to the form $K(t)=\lambda(t)^{L}$, with the extracted $\lambda(t)$ plotted as black lines. The excellent agreement between the largest eigenvalues and $\lambda(t)$ demonstrates that, within the accessible pre-ramp time window, the SFF is indeed governed by the non-degenerate leading eigenvalues of the transfer matrix. This behavior contrasts with Floquet models such as the kicked Ising model [chanSolutionMinimalModel2018c], where the SFF is controlled by a highly degenerate set of leading eigenvalues from the outset.
Within the accessible pre-ramp time window, the dynamics are governed by the non-degenerate leading eigenvalue, which is size independent, cannot explain the size-dependent onset of the RMT ramp (the Thouless time, black dashed in Fig. 3(a)). Understanding how the system crosses over from the leading-eigenvalue regime to the ramp regime thus requires access to longer times.
At present, a direct evaluation of the eigenvalues in the ramp region is hindered by rapid bond growth in Krylov iteration and the lack of convergence in non-Hermitian DMRG. Looking beyond the numerically accessible regime, the observed broadening of the confidence intervals and the near-degeneracy of the leading eigenvalues suggest a possible crossover scenario. In particular, we conjecture that around the Thouless time the leading eigenvalues become increasingly near-degenerate, such that at finite sizes multiple eigenvalues contribute comparably, producing the ramp, whereas in the thermodynamic limit the single largest eigenvalue eventually dominates. This conjectured scenario is consistent with the size dependence of the Thouless time and is further illustrated by a toy-model analysis [supplementaryMaterial].
V Summary and outlook
We propose SeTN, a tensor network approach for studying chaotic dynamics in general disordered quantum systems. By encoding disorder into an auxiliary layer and performing the disorder averaging separately, SeTN restores spatial translational invariance and enables accurate simulations directly in the thermodynamic limit through a compact MPO transfer matrix.From the singular value spectrum, we derive and numerically confirm a universal criterion for efficient disorder encoding, $n\gg\alpha^{2}t^{2}$, which indicates that encoding is more efficient under weak disorder, a regime that is typically chaotic and challenging for conventional TN methods, making disorder-averaged transfer matrices numerically feasible for generic Hamiltonians.
We compute the leading eigenvalue of the SeTN-derived transfer matrix for the SFF in the disordered TFIM, and show that it accurately captures the scaling behavior of the SFF within the accessible time range. Based on the behavior observed in this regime, we conjecture that the crossover from single-eigenvalue dominance to random-matrix-theory behavior is associated with an increasing near-degeneracy among the leading transfer-matrix eigenvalues. Clarifying how this near-degeneracy develops and how it controls the emergence and system-size dependence of the Thouless time remains an important open problem for future work.
SeTN is broadly applicable to observables with multilayer TN representations, including Rényi entropies [chanSolutionMinimalModel2018c, zhouEmergentStatisticalMechanics2019], out-of-time-ordered correlators (OTOCs) [nahumOperatorSpreadingRandom2018c, vonkeyserlingkOperatorHydrodynamicsOTOCs2018a, chanSolutionMinimalModel2018c], and spin-glass order parameters [huseLocalizationprotectedQuantumOrder2013a, pekkerHilbertGlassTransitionNew2014a], offering a powerful platform for investigating the impact of disorder on dynamical phenomena such as diffusion and hydrodynamics [khemaniOperatorSpreadingEmergence2018c, rakovszkyDiffusiveHydrodynamicsOutofTimeOrdered2018c]. SeTN also serves as a bridge between local averaging schemes and other concepts, such as Feynman-trajectory picture [vonkeyserlingkOperatorHydrodynamicsOTOCs2018a, zhouEmergentStatisticalMechanics2019, zhouEntanglementMembraneChaotic2020], influence functionals [nguyenCorrelationFunctionsTensor2024, sonnerInfluenceFunctionalManybody2021a, yeConstructingTensorNetwork2021, leroseInfluenceMatrixApproach2021b] and entanglement barriers [rathEntanglementBarrierIts2023]. Overall, our construction provides a unified framework for exploring the interplay of locality, unitarity, disorder, and chaos in static many-body systems. In this sense, SeTN complements recent developments such as emergent symmetries in disorder-averaged dynamics [erpeldingExploitingEmergentSymmetries2025a] and translationally invariant MPO descriptions of thermal ensembles [vervoortExtractingAverageProperties2025], and opens multiple avenues for future investigation.
Acknowledgment.— This work was supported in part by National Natural Science Foundation of China Grants No. 12474353 and Beijing Natural Science Foundation (Grant No. 1232025). SJR acknowledges supports from Peng Huanwu Visiting Professor Program, Chinese Academy of Sciences, and Academy for Multidisciplinary Studies, Capital Normal University. D.W. acknowledges support from the National Research Foundation, Singapore and the Agency for Science, Technology and Research (A*STAR) under the Quantum Engineering Programme (NRF2021-QEP2-02-P03) and from CQT Core Funding Grant CQT_SUTD_2025_01.
Data availability.— All data are available from the authors upon reasonable request.